It is correct to conclude that the:

a) The monetary base is equal to cash in circulation.

b) The money supply is equal to the monetary base.

c) The money supply does not include sight deposits.

d) The monetary base is equal to banks’ cash reserves.

e) None of the above



It is incorrect to conclude that:

a) When the income tax rate is 0 and the multiplier is 10, the MPC = 0.9.

b) If the income tax rate is 0.1 and the MPC = 1, then the multiplier is equal to 10. c) If the income tax rate is 0.2 and the MPC = 0.75, then the multiplier is equal to 0.9375.

d) If the tax rate is 1 the MPC cannot be calculated.

e) None of the above


Suppose the consumption function is C = 100 + 0.95YD. If the tax rate changes from t = 0 to t = 30%, then the increase in government spending that leaves the equilibrium income unaffected is:

a) 1,478

b) 2,000

c) 2,570

d) 10,280

e) 570


1
Expert's answer
2021-12-10T14:31:38-0500

d) The monetary base is equal to banks’ cash reserves.

d) If the tax rate is 1 the MPC cannot be calculated.

If t = 0, then T = 0, so C = 100 + 0.95(Y - T) = 100 + 0.95Y.

If t = 0.3, then T = t"\times" Y = 0.3Y, so C = 100 + 0.95(Y - 0.3Y) = 100 + 0.665Y

Y=570

e) 570